You download a small HTML file that references 8 small objects (that is, ignore transmission times). All these 8 objects reside on the same server. One RTT is 300 ms. How much time does it take to download all 9 objects using nonpersistent but concurrent (max 6) HTTP connections? Give answer in seconds to one decimal digit.
Answer #1

Non-persistent connection required 2RTTs per object.

Here 6 HTTP connections can be concurrent so for total 9 objects it requires = 2*(9/6)RTTs = 4RTTs

So total time = 2RTTs + 4 RTTs = 6RTTs = 6*300ms = 1.8 second